Company XYZ closed at ​$52.29 per share with a​ P/E ratio of 14.03.

A. How much were earnings per​ share?

B .Does the stock seem​ overpriced, underpriced, or about right given that the historical​ P/E ratio is​ 12-14?
Answer #1

Given

a)

P/E ratio = 14.03

P/E ratio=price per share/Earnings per share

Here

Price per share = $52.29

Earnings per share = 52.29/14.03

Earnings per share = $3.727

b)

historical​ P/E ratio is​ 12-14

P/E ratio is greater than average historical P/E ratio.

Stock seems overpriced